# Quantum Energy Teleportation: A Novel Protocol Relaxing Previous Constraints

### v1

@@ VI Acknowledgments
The details of the three IBM quantum hardware backends are provided in this section. The qubit distribution maps for ibm_brussels and ibm_kyiv are shown in Fig. 3, while the map for ibm_torino is presented in Fig. 4. In all three quantum circuits, qubit 0 is used for Alice’s qubit, and qubit 1 is used for Bob’s qubit.
@@
Table 3 provides detailed information about the three IBM quantum hardware backends, including the properties of the qubits used and the properties of the single-qubit and two-qubit gates employed in the experiments. This information was retrieved from IBM Quantum (2025) on January 13, 2025, the day when the quantum circuits were executed on these backends.
@@
Figure 3: Qubit distribution map for the backends of ibm_brussels and ibm_kyiv is demonstrated. Qubit 0 is used for Alice’s qubit, and qubit 1 is used for Bob’s qubit.
@@
Figure 4: Qubit distribution map for the backend of ibm_torino is demonstrated. Qubit 0 is used for Alice’s qubit, and qubit 1 is used for Bob’s qubit.

### v2

@@ Appendix C Details of the Quantum Hardware
The details of the three IBM quantum hardware backends are provided in this section. The qubit distribution maps for ibm_brussels and ibm_kyiv are shown in Figure 4, while the map for ibm_torino is presented in Figure 5. In all three quantum circuits, qubit 0 is used for Alice’s qubit, and qubit 1 is used for Bob’s qubit.
@@
Table 3 provides detailed information about the three IBM quantum hardware backends, including the properties of the qubits used and the properties of the single-qubit and two-qubit gates employed in the experiments. This information was retrieved from IBM Quantum (2025) on 13 January 2025, the day when the quantum circuits were executed on these backends.
@@
Figure 4: The qubit distribution map for the backends of ibm_brussels and ibm_kyiv is demonstrated. Qubit 0 is used for Alice’s qubit, and qubit 1 is used for Bob’s qubit.
@@
Figure 5: The qubit distribution map for the backend of ibm_torino is demonstrated. Qubit 0 is used for Alice’s qubit, and qubit 1 is used for Bob’s qubit.